OpenStudy (anonymous):

Input the equation for the following line. Reduce all fractional answers to lowest terms. A line that passes midway between and is parallel to both x - y = 6 and x - y = -6.

OpenStudy (anonymous):

First, put both into slope-intercept form: -y=-x+6-->y+x-6 and -y=-x-6-->y=x+6 Parallel lines have the same slope. In this case, the slope is 1, so that is going to be the slope of our third line. What is midway between -6 and 6? Well, 0. So plug our slope and vertical (0) into the new equation and we get: y=x Put that into standard form (because the other equations are) and we get: x-y=0